# Tropicalizing Binary Geometries

### v1

@@ 5 The complex of axially symmetric phylogenetic trees
Furthermore, for every diagonal $d$ of $P$ we have a diagonal $d^{\prime}$ that is symmetric to $d$ with respect to $l$. The two diagonals $d$ and $d^{\prime}$ coincide if and only if $d$ is $l$ or is one of the $n-1$ diagonals perpendicular to $l$. This allows us to consider axially symmetric subdivisions of $P$: those which contain a diagonal if and only if they also contain the one symmetric to it. Note that such a subdivision may not contain diagonals that cross $l$ but are not perpendicular to $l$.
@@
Now we can give an analogue of Definition 2.3. Note that since an ASPT is compatible with multiple ASDOs, the various $\Delta_{\mathrm{as}}(\alpha)$ share vertices.
@@
The complex of axially symmetric phylogenetic trees $\Theta_{\mathrm{as}}(n)$ is the union of $\Delta_{\mathrm{as}}(\alpha)$ over all ASDOs $\alpha$. The vertex set of the simplicial complex $\Theta_{\mathrm{as}}(n)$ is the set of all minimal ASPTs. A subset of vertices forms a face if and only if it forms a face in one of the $\Delta_{\mathrm{as}}(\alpha)$.
@@
For a non-leaf edge $e$ of $T$ we may again consider the phylogenetic tree obtained from $(T,\varphi)$ by contracting $e$ but the result might not be axially symmetric. However, if we contract both $e$ and $\iota(e)$ (or just $e$ if it is symmetric to itself), then what we obtain is axially symmetric. We call such an operation a symmetric edge contraction. Since it corresponds to the deletion of a pair of mutually symmetric diagonals from the subdivision (or of a single diagonal symmetric to itself), we have the following.
@@ 6 The (sub)complex of centrally symmetric phylogenetic trees
A dihedral ordering $\alpha$ of $N$ is centrally symmetric if it is given by an ordering $(\sigma_{1},\dots,\sigma_{2n})$ of $N$ satisfying $\sigma_{i+n}=-\sigma_{i}$ for all $i\in[n]$ (hence, any other ordering representing $\alpha$ also has this property). There are $2^{n-2}(n-1)!$ centrally symmetric dihedral ordering (or CSDOs).
Let $P$ be a regular $2n$-gon with edges labeled according to a CSDO $\alpha$. A subdivision of $P$ is centrally symmetric if with every diagonal it also includes its reflection in the center of $P$. By considering the phylogenetic trees corresponding to such subdivisions we obtain a notion of centrally symmetric phylogenetic trees (CSPTs) and a notion of compatibility between CSPTs and CSDOs.
There are $n(n-1)$ coarsest nontrivial centrally symmetric subdivisions of $P$: the $n$ formed by one longest diagonal and $n(n-2)$ more formed by a pair of distinct diagonals that are centrally symmetric to each other. Again, we say that a CSPT compatible with $\alpha$ is minimal if it corresponds to a coarsest nontrivial centrally symmetric subdivision of $P$. Such trees also have two or three non-leaf vertices.
@@
The faces of $\Delta_{\mathrm{cs}}(\alpha)$ are in bijection with centrally symmetric subdivisions of $P$ or, in other words, with CSPTs compatible with $\alpha$. This is a flag complex of pure dimension $n-2$; we refer to $\Delta_{\mathrm{cs}}(\alpha)$ as the dual cyclohedron since its face poset is dual to that of a cyclohedron. Again we consider the respective union.
@@
The complex of centrally symmetric phylogenetic trees $\Theta_{\mathrm{cs}}(n)$ is the union of $\Delta_{\mathrm{cs}}(\alpha)$ over all CSDOs $\alpha$. This is the simplicial complex whose vertex set consists of all minimal CSPTs with a subset forming a face if and only if it forms a face in one of the $\Delta_{\mathrm{cs}}(\alpha)$.
Similarly to the previous cases, the faces of $\Theta_{\mathrm{cs}}(n)$ are in bijection with the set of all CSPTs. An important observation is that CSPTs are also axially symmetric.
@@
In Figure A the subcomplex $\Theta_{\mathrm{cs}}(3)$ is highlighted in bold. It has 10 vertices and 12 edges. One may check that the respective trees are indeed centrally symmetric. The subcomplex $\Delta_{\mathrm{cs}}(\alpha)$ for $\alpha$ given by $(1,-2,3,-1,2,-3)$ is highlighted in blue; it is the dual-cyclohedron, i.e. a 6-cycle.
@@ 7 $\mathcal{M}_{\mathrm{C}_{n-1}}$ and its tropicalization
In this section we introduce the variety $\mathcal{M}_{\mathrm{C}_{n-1}}$ as the (very affine part of the) binary geometry defined by the dual cyclohedron. We then give a conjectural description of $\operatorname{Trop}\mathcal{M}_{\mathrm{C}_{n-1}}$ as of the space of ASPTs. We also conjecture that the signed tropicalizations of $\mathcal{M}_{\mathrm{C}_{n-1}}$ are the subfans corresponding to the dual-associahedron and the dual-cyclohedron subcomplexes.
Let $n$ and $N$ be as in the previous section and let $\alpha_{0}$ denote the CSDO of $N$ given by the ordering $(1,\dots,n,-1,\dots,-n)$. We set $\Delta_{\mathrm{cs}}(n)=\Delta_{\mathrm{cs}}(\alpha_{0})$. As in the case of $\mathcal{M}_{0,n}$ we enumerate the vertices of $\Delta_{\mathrm{cs}}(n)$ by certain pairs of elements of $N$. Recall that these vertices correspond to minimal CSPTs compatible with $\alpha_{0}$ and, thus, to coarsest nontrivial centrally symmetric subdivisions of the regular $2n$-gon $P$ with edges labeled in accordance with $\alpha_{0}$.
@@
Our description of $\operatorname{Trop}\mathcal{M}_{\mathrm{C}_{n-1}}=\operatorname{Trop}I_{n}$ will be given in terms of pseudometrics on symmetric phylogenetic trees. Consider an ASPT $(T,\varphi)$, and let $E$ denote the set of non-leaf edges of $T$ and $\iota$ denote the symmetry of $(T,\varphi)$. A symmetric realization of $(T,\varphi)$ is an assignment of lengths to the edges in $E$ given by a vector $l=(l_{e})_{e\in E}\in\mathbb{R}_{\geq 0}^{E}$ such that $l_{e}=l_{\iota(e)}$ for any $e\in E$. Such a realization determines distances between pairs of vertices of $T$ where the lengths of leaf edges are assumed to be zero. For $i,j\in N$ we denote by $d_{(T,\varphi),l}(i,j)$ the resulting distance function between leaf vertices $\varphi(i)$ and $\varphi(j)$. Evidently, this function satisfies $d(i,j)=d(-i,-j)$.
Let $\operatorname{SR}(T,\varphi)$ denote the of all symmetric realizations of $(T,\varphi)$. We associate a cone in $\mathbb{R}^{D}$ with the tree $(T,\varphi)$ as follows:
@@
Our second conjecture concerns the sign patterns occurring in $\mathcal{M}_{\mathrm{C}_{n-1}}(\mathbb{R})$ and the corresponding signed tropicalizations. In the following note that a dihedral ordering of $N$ cannot be both axially symmetric and centrally symmetric, hence there is a total of $2^{n-2}(n+1)(n-1)!$ dihedral orderings that have one of these properties.
@@
This is an extension of a conjecture made in [2].
Conjecture 7.7 ([2, Subsection 11.3]).
The number of sign patterns occurring in $\mathcal{M}_{\mathrm{C}_{n-1}}(\mathbb{R})$ is $2^{n-2}(n+1)(n-1)!$.
The case of the positive tropicalization, i.e. $\tau=(1,\dots,1)$, is also studied in loc. cit. In this case the we expect the proposed bijection to give $\tau=\tau_{\alpha_{0}}$ which would agree with the following result.
Theorem 7.8 ([2, Theorem 9.2]).
The positive tropicalization $\operatorname{Trop}_{>0}\mathcal{M}_{\mathrm{C}_{n-1}}$ is combinatorially equivalent to a fan over the dual-cyclohedron complex $\Delta_{\mathrm{cs}}(n)$.
@@ 8 Example: $\operatorname{Trop}\mathcal{M}_{\mathrm{C}_{2}}$
Acknowledgments. We thank Thomas Lam for introducing us to these questions and giving insightful suggestions. We also thank Karin Baur and Zachary Greenberg for helpful discussions.

### v2

@@ 5 The complex of axially symmetric phylogenetic trees
Furthermore, for every diagonal $d$ of $P$ we have a diagonal $d^{\prime}$ that is symmetric to $d$ with respect to $l$. The two diagonals $d$ and $d^{\prime}$ coincide if and only if $d$ is $l$ or is one of the $n-1$ diagonals perpendicular to $l$. This allows us to consider axially symmetric subdivisions of $P$: subdivisions which contain a diagonal if and only if they also contain the one symmetric to it. Such a subdivision may not contain diagonals that cross $l$ but are not perpendicular to $l$.
@@
Now we can give an analogue of Definition 2.3. Note that, similarly to the phylogenetic trees considered in Section 2, any ASPT is compatible with multiple ASDOs. Hence, the union of the various $\Delta_{\mathrm{as}}(\alpha)$ is not disjoint.
@@
The complex of axially symmetric phylogenetic trees $\Theta_{\mathrm{as}}(n)$ is the union of $\Delta_{\mathrm{as}}(\alpha)$ over all ASDOs $\alpha$, with vertices represented by the same phylogenetic trees identified. The vertex set of the simplicial complex $\Theta_{\mathrm{as}}(n)$ is the set of all minimal ASPTs. A subset of vertices forms a face if and only if it forms a face in one of the $\Delta_{\mathrm{as}}(\alpha)$.
@@
For a non-leaf edge $e$ of $T$ we may again consider the phylogenetic tree obtained from $(T,\varphi)$ by contracting $e$, but the result might not be axially symmetric. However, if we contract both $e$ and $\iota(e)$ (or just $e$ if it is symmetric to itself), then what we obtain is axially symmetric. We call such an operation a symmetric edge contraction. Since it corresponds to the deletion of a pair of mutually symmetric diagonals from the subdivision (or of a single diagonal symmetric to itself), we have the following.
@@ 6 The (sub)complex of centrally symmetric phylogenetic trees
A dihedral ordering $\alpha$ of $N$ is centrally symmetric if it is given by an ordering $(\sigma_{1},\dots,\sigma_{2n})$ of $N$ satisfying $\sigma_{i+n}=-\sigma_{i}$ for all $i\in[n]$ (hence, any other ordering representing $\alpha$ also has this property). There are $2^{n-2}(n-1)!$ centrally symmetric dihedral orderings (or CSDOs).
Let $P$ be a regular $2n$-gon with edges labeled according to a CSDO $\alpha$. A subdivision of $P$ is centrally symmetric if with every diagonal $d$ it also includes the diagonal centrally symmetric to $d$. By considering phylogenetic trees arising from such subdivisions, we obtain a notion of centrally symmetric phylogenetic trees (CSPTs) and a notion of compatibility between CSPTs and CSDOs.
There are $n(n-1)$ coarsest nontrivial centrally symmetric subdivisions of $P$: the $n$ formed by one longest diagonal and $n(n-2)$ more formed by a pair of distinct diagonals that are centrally symmetric to each other. Again, we say that a CSPT compatible with $\alpha$ is minimal if it arises from a coarsest nontrivial centrally symmetric subdivision of $P$. Such trees also have two or three non-leaf vertices.
@@
The faces of $\Delta_{\mathrm{cs}}(\alpha)$ are in bijection with centrally symmetric subdivisions of $P$ or, in other words, with CSPTs compatible with $\alpha$. This is a flag complex of pure dimension $n-2$; we refer to $\Delta_{\mathrm{cs}}(\alpha)$ as a dual cyclohedron since its face poset is dual to that of a cyclohedron. Again, a CSPT will be compatible with multiple CSDOs, making the union of the $\Delta_{\mathrm{cs}}(\alpha)$ non-disjoint.
@@
The complex of centrally symmetric phylogenetic trees $\Theta_{\mathrm{cs}}(n)$ is the union of $\Delta_{\mathrm{cs}}(\alpha)$ over all CSDOs $\alpha$, with vertices represented by the same phylogenetic trees identified. This is the simplicial complex whose vertex set consists of all minimal CSPTs with a subset forming a face if and only if it forms a face in one of the $\Delta_{\mathrm{cs}}(\alpha)$.
Similarly to the previous cases, the faces of $\Theta_{\mathrm{cs}}(n)$ are in bijection with CSPTs. An important observation is that CSPTs are also axially symmetric.
@@
In Figure A the subcomplex $\Theta_{\mathrm{cs}}(3)$ is highlighted in bold. It has 10 vertices and 12 edges. One may check that the respective trees are indeed centrally symmetric. The subcomplex $\Delta_{\mathrm{cs}}(\alpha)$ for $\alpha$ given by $(1,-2,3,-1,2,-3)$ is highlighted in blue; it is the dual cyclohedron, i.e. a 6-cycle.
@@ 7 $\mathcal{M}_{\mathrm{C}_{n-1}}$ and its tropicalization
In this section we first recall the definition of the type C cluster configuration space $\mathcal{M}_{\mathrm{C}_{n-1}}$. We follow [2] where $\mathcal{M}_{\mathrm{C}_{n-1}}$ is realized as the (very affine part of the) binary geometry defined by the dual cyclohedron. We then give a conjectural description of $\operatorname{Trop}\mathcal{M}_{\mathrm{C}_{n-1}}$ as of the space of ASPTs. We also conjecture that the signed tropicalizations of $\mathcal{M}_{\mathrm{C}_{n-1}}$ are the subfans corresponding to the dual-associahedron and the dual-cyclohedron subcomplexes.
Let $n$ and $N$ be as in the previous section and let $\alpha_{0}$ denote the CSDO of $N$ given by the ordering $(1,\dots,n,-1,\dots,-n)$. We set $\Delta_{\mathrm{cs}}(n)=\Delta_{\mathrm{cs}}(\alpha_{0})$. As in the case of $\mathcal{M}_{0,n}$, we enumerate the vertices of $\Delta_{\mathrm{cs}}(n)$ by certain pairs of elements of $N$. Recall that these vertices correspond to minimal CSPTs compatible with $\alpha_{0}$ and, thus, to coarsest nontrivial centrally symmetric subdivisions of the regular $2n$-gon $P$ with edges labeled in accordance with $\alpha_{0}$.
@@
Our conjectured description of $\operatorname{Trop}\mathcal{M}_{\mathrm{C}_{n-1}}=\operatorname{Trop}I_{n}$ will be given in terms of pseudometrics on symmetric phylogenetic trees. Consider an ASPT $(T,\varphi)$, and let $E$ denote the set of non-leaf edges of $T$ and $\iota$ denote the symmetry of $(T,\varphi)$. A symmetric realization of $(T,\varphi)$ is an assignment of lengths to the edges in $E$ given by a vector $l=(l_{e})_{e\in E}\in\mathbb{R}_{\geq 0}^{E}$ such that $l_{e}=l_{\iota(e)}$ for any $e\in E$. Such a realization determines distances between pairs of vertices of $T$ where the lengths of leaf edges are assumed to be zero. For $i,j\in N$ we denote by $d_{(T,\varphi),l}(i,j)$ the resulting distance function between leaf vertices $\varphi(i)$ and $\varphi(j)$. Evidently, this function satisfies $d(i,j)=d(-i,-j)$.
Let $\operatorname{SR}(T,\varphi)$ denote the set of all symmetric realizations of $(T,\varphi)$. We associate a cone in $\mathbb{R}^{D}$ with the tree $(T,\varphi)$ as follows:
@@
Our second conjecture concerns the sign patterns occurring in $\mathcal{M}_{\mathrm{C}_{n-1}}(\mathbb{R})$ and the corresponding signed tropicalizations. In the following note that a dihedral ordering of $N$ cannot be both axially symmetric and centrally symmetric, thus there are $2^{n-2}(n+1)(n-1)!$ dihedral orderings with one of these properties.
@@
The conjecture above extends a conjecture made in [2, Subsection 11.3] where the authors, based on computational data, hypothesize that the number of sign patterns occurring in $\mathcal{M}_{\mathrm{C}_{n-1}}(\mathbb{R})$ is $2^{n-2}(n+1)(n-1)!$.
The case of the positive tropicalization, i.e. of the sign pattern $(1,\dots,1)$, is also studied in loc. cit. We expect the proposed bijection to give $\tau_{\alpha_{0}}=(1,\dots,1)$. This would agree with [2, Theorem 9.2] which implies that $\operatorname{Trop}_{>0}\mathcal{M}_{\mathrm{C}_{n-1}}$ is combinatorially equivalent to a fan over the dual-cyclohedron complex $\Delta_{\mathrm{cs}}(n)$.
@@ 8 Example: $\operatorname{Trop}\mathcal{M}_{\mathrm{C}_{2}}$
Acknowledgments. We thank Thomas Lam for introducing us to these topics and for his insightful suggestions. Thanks also to Karin Baur, Zachary Greenberg and Giulio Salvatori for helpful discussions, and to Chiara Meroni and Hannah Tillman-Morris for their feedback on the paper.